Sandy is 35 years older than Tony the sum of their ages is 43. How old are they?
Answer by ikleyn(52776)


Let  x  be the Tony's age in years.
Then Sandy's age is  x + 35,  according to the condition.
Next,  you have an equation

(x + 35) + x = 43.

Simplify and solve it:
2x + 35 = 43,
2x = 43 - 35,
2x = 8,
x = 4.

Answer. Tony is  4  years old,  Sandy is  4 + 35 = 39 years old.